The 100th term of an AP is 98, and the common difference is 2. Find the first three terms.
Answer by richwmiller(17219)

an = a + (n - 1)d
98 = a1 + 2*99
98=198+a1
100+a1=0
a1=-100
a2=-98
a3=-96
BTW
a51=-100+2*50
a51=0